OpenStudy (anonymous):

Write each statement as an equation, and find the number. Ten more than one-third of a number is the same as the number decreased by two I Need to show all the work. I Am Horrible At Math Can Someone Please Take Me Through The Steps to get positive 18For The Answer

OpenStudy (anonymous):

let the number Ten more than one third of a number is the same (equal to) as the number decreased by 2 in the form of an equation it will look like this: 10 + (1/3)x = x - 2

OpenStudy (anonymous):

let the number equal x by the way. I meant to write that.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Now you are left with an algebraic expression, for which you can solve for x

OpenStudy (anonymous):

I am Horrible at solving with fractions please help

OpenStudy (anonymous):

you get 12+1/3x=x Right?

OpenStudy (anonymous):

yes that is right

OpenStudy (anonymous):

Then I would subtract 1/3 from both sides correct

OpenStudy (anonymous):

now remember that x is the same as (3x/3), so you can use that to combine the fractions easier

OpenStudy (anonymous):

I would write it as 12 = (3x/3) - (1x/3) 12 = 2x/3 multiply the 12 by 3 then divide by 2 x = 18